5. Most recent financial statement is shown

Income Statement Sales Costs Taxable Balance Sheet $18,900 Current $ 11,700 Debt $15,700 assets 12,800 Fixed assets 26,500 Equity 22,500 $ 6,100 Tota $38,200 $38,200 income Taxes (21%) 1,281 Net income $ 4,819 Assets and costs are proportional to sales. Debt and equity are not. The company maintains a constant 30 percent dividend payout ratio. What is the sustainable growth rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.
Answer #1

As growth rate = return on equity * plowback ratio(ROE * b),

Sustainable growth rate = (ROE × b) / [1 – (ROE × b)]

ROE = Net income / Equity

=4819/22500

=21.418%

plowback ratio = 1- dividend payout ratio = 1- 0.3 = 0.7

Sustainable growth rate = (ROE × b) / [1 – (ROE × b)]

=(21.418% * 0.7) / 1- (21.418% * 0.7)

=17.64%